### v1

# ShortcutsBench: A Large-Scale Real-world Benchmark for API-based Agents

Abstract
Recent advancements in integrating large language models (LLMs) with application programming interfaces (APIs) have gained significant interest in both academia and industry. These API-based agents, leveraging the strong autonomy and planning capabilities of LLMs, can efficiently solve problems requiring multi-step actions. However, their ability to handle multi-dimensional difficulty levels, diverse task types, and real-world demands through APIs remains unknown. In this paper, we introduce ShortcutsBench, a large-scale benchmark for the comprehensive evaluation of API-based agents in solving tasks with varying levels of difficulty, diverse task types, and real-world demands. ShortcutsBench includes a wealth of real APIs from Apple Inc.’s operating systems, refined user queries from shortcuts, human-annotated high-quality action sequences from shortcut developers, and accurate parameter filling values about primitive parameter types, enum parameter types, outputs from previous actions, and parameters that need to request necessary information from the system or user. Our extensive evaluation of agents built with $5$ leading open-source (size >= 57B) and $4$ closed-source LLMs (e.g. Gemini-1.5-Pro and GPT-3.5) reveals significant limitations in handling complex queries related to API selection, parameter filling, and requesting necessary information from systems and users. These findings highlight the challenges that API-based agents face in effectively fulfilling real and complex user queries. All datasets, code, and experimental results will be available at https://github.com/eachsheep/shortcutsbench.
1 Introduction
Large language model based agents (LLM-based agents) (Wang et al., 2023b; Xi et al., 2023) built on application programming interfaces (APIs) (Qin et al., 2023; Huang et al., 2023) have recently gained significant interest in both academia (Shen et al., 2024; Wang et al., 2023b) and industry (Microsoft, 2024; OpenAI, 2024c). By integrating LLM with APIs, these agents can access real-time information (OpenAI, 2024a; Microsoft, 2024), reduce hallucination with external knowledge (Li et al., 2023a; Gao et al., 2023), and automatically plan and complete complex tasks that need multi-step actions (Gravitas, 2024; Pan et al., 2023). Many of these agents (OpenAI, 2024c; Microsoft, 2024; Gravitas, 2024) have also demonstrated commendable performance on simple tasks involving only a few actions such as “Check the weather ① and tell me ②” (OpenAI, 2024c). These impressive performances raise an important question: Do these API-based agents truly possess the capability to generate complex action sequences for real demands with real APIs?
Some existing benchmarks/datasets (Huang et al., 2023; Qin et al., 2023; Patil et al., 2023; Tang et al., 2023; Li et al., 2023b; Xu et al., 2023; Zhuang et al., 2024; Schick et al., 2024; Hao et al., 2024) have attempted to evaluate API-based agents. However, they have three limitations: First, the APIs (a.k. tools available to the agent) lack richness, and the queries (a.k. the task to the agent) lack complexity (Table 1). They either involve a limited number of APIs (Li et al., 2023b; Xu et al., 2023; Zhuang et al., 2024) or cover only a few numbers of apps (Patil et al., 2023; Tang et al., 2023; Schick et al., 2024; Hao et al., 2024) (an app may have one or multiple APIs), or the difficulties of the queries are limited in a narrow range, with the average action length ranges from $1$ to $5.9$ (Table  1). This lack of richness and complexity makes it difficult to effectively distinguish the capabilities of different agents, particularly on larger and smarter LLMs like Gemini-1.5 Pro DeepMind (2024) and GPT-3.5-turbo OpenAI (2023a). Second, the APIs lack realism as they may be manually crafted, and the queries fail to reflect actual user demands since they mey be either created by hand or generated directly by ChatGPT OpenAI (2023a) without verifying real user demands (Table 1). Moreover, they only cover the evaluation of API selection, lacking a study on API parameter filling (Table 1). Efficient and accurate parameter filling is essential for an agent to complete tasks successfully. Thirdly, they don’t adequately evaluate the agent’s ability to ask systems or the users for the necessary information to resolve the missing information for solving the queries (Table 1). This is crucial as a user’s query may be implicit or may not provide all the information an agent needs to solve the task (Qin et al., 2023; Qian et al., 2024) effectively.
In this paper, we present ShortcutsBench. To our best known, ShortcutsBench is the first large-scale real API-based agent benchmark considering APIs, queries, and corresponding action sequences. ShortcutsBench provides rich real APIs, queries with various difficulties and task types, high-quality human-annotated action sequences by shortcuts developers, along queries from real user demands. Moreover, it also provides precise values for parameter filling, including primitive data types, enum types, and the use of output from previous actions for parameter values, as well as evaluations of the agent’s awareness in requesting necessary information from the system or user. Furthermore, the scale of APIs, queries, and the corresponding action sequences in ShortcutsBench is comparable or even better to benchmarks and datasets created by LLM or modified by existing datasets. The overall comparison between ShortcutsBench and existing benchmarks / datasets is listed in Table 1.
We conducted extensive evaluations of API-based agents from $9$ leading LLMs on  ShortcutsBench, including the evaluation of API selection, parameter value filling, and recognition of the need for input from the system or the user. The chosen LLMs including four closed-sourced LLMs like Gemini-1.5-Pro (DeepMind, 2024) and GPT-3.5-turbo (OpenAI, 2024b), and five open-source LLMs like LLaMA-3-70B (Meta, 2024) and QWen-2-70B (Qwen, 2024b). Our findings highlight the limitations of these agents in addressing real, rich, and complex user queries. In summary, this paper makes the following key contributions:
  • To our best known, we have built the most realistic, rich, and comprehensive API-based agent benchmark. This benchmark is even comparable in scale to existing benchmarks/datasets built using LLMs (Table 1).
  • We evaluated nine most advanced and mainstream LLM-based agents on all operations required to complete user queries, including API selection, parameter filling, and their awareness to request necessary information from the system or user when needed.
  • We obtained massive interesting conclusions such as (1) The performance gap between open-source and closed-source LLMs has become very small; (2) Existing LLMs still have significant shortcomings in multi-step reasoning; (3) Extracting necessary parameters from queries is the most challenging task in parameter filling; (4) There is a substantial lack of awareness in agents when it comes to requesting necessary information.
Table 1: ShortcutsBench has a great advantage in the realness, richness, and complexity of APIs, queries, and corresponding action sequences, the validity of action sequences, accurate parameter value filling, the awareness for asking information from the system or the users, and the overall scale.
Resource
Shortcuts
Bench
Meta
Tool
Tool
LLM
API
Bench
Tool
Alpaca
API
Bank
Tool
Bench
Tool
QA
2023
2023
2023
2023
2023b
2023
2024
Real API?
Demand-driven Query?
Human-Annotated Act.?
Multi-APIs Query?
Multi-Step Act.?
Prec. Val. for Para. Fill?
Awareness for Ask Info?
# Apps 88 390 3451 3 53 400 8 13
# APIs 1414 390 16464 1645 53 400 232 13
# Queries 7628 21127 12657 17002 3938 274 2726 1530
# Avg APIs 7.86 1.0 2.3 1.0 1.0 2.1 5.4 13.0
# Avg Actions 21.46 1.0 4.0 1.0 1.0 2.2 5.9 1.0
2 Related Work
API-based agents.
API-based agents treat APIs as tools (Huang et al., 2023; Qin et al., 2023; Patil et al., 2023; Tang et al., 2023; Li et al., 2023b; Xu et al., 2023; Zhuang et al., 2024; Schick et al., 2024; Hao et al., 2024; Zhu et al., 2023; Gravitas, 2024; AgentGPT, 2023). They accept queries, generate action sequences based on queries and provided APIs, and generate next action depends on the history actions (Yao et al., 2022). Related work about API-based agents can generally be categorized into $3$ types depending on the objective: (1) Task-specific enhancement focuses on improving the agent’s ability to solve a specific type of task like game and question-answering (Hao et al., 2024; Zhu et al., 2023; Gravitas, 2024; AgentGPT, 2023). (2) Data-driven workflows emphasize the importance of data by researching how to construct workflows to get action sequences, enabling generated data to fine-tune the model (Qin et al., 2023; Patil et al., 2023; Tang et al., 2023; Xu et al., 2023; Zhuang et al., 2024; Schick et al., 2024). (3) Agent evaluation studies the assessment of agents (Huang et al., 2023; Li et al., 2023b).
Code-based agents.
Code-based agents use code generated as a medium for interaction with the external environment (OpenAI, 2023b; Wang et al., 2023a; OpenInterpreter, 2024; Wu et al., 2023). They accept queries, generate scripts in programming languages such as Python (OpenAI, 2023b), JavaScript (Wang et al., 2023a), or Shell (OpenInterpreter, 2024), and then input the code into interpreters. The results of the interpreter are then returned to the agent, which is used to help determine the next action in code generation. Currently, these approaches primarily focus on enhancing agent performance in specific tasks by incorporating additional knowledge (Wang et al., 2023a; Wu et al., 2023), increasing feedback (OpenAI, 2023b; Wang et al., 2023a; OpenInterpreter, 2024; Wu et al., 2023), and decomposing tasks (OpenAI, 2023b; Wang et al., 2023a; OpenInterpreter, 2024; Wu et al., 2023).
Digital Automation Platforms (DAPs).
DAPs (Abdou et al., 2021; Coze, 2023; Rahmati et al., 2017; Chakraborti et al., 2020) refer to software tools or services designed to optimize workflows through automation. DAPs leverage technologies such as robotic process automation (RPA) (Chakraborti et al., 2020) and low-code/no-code development tools to achieve the goals. DAPs like Zapier (Abdou et al., 2021; Rahmati et al., 2017), Make (Make, 2023), and IFTTT (Abdou et al., 2021; Rahmati et al., 2017) offer extensive APIs that enable users to create automated workflows. Similarly, DAPs such as Microsoft Power Automate (Abdou et al., 2021) and Tasker (Dias, 2024) are primarily used to build workflows on Azure and Android, respectively. Recently, with the rise of LLM-based agents, platforms like Coze (Coze, 2023) and Dify (Dify, 2023) have emerged as “agent construction platforms”. Functionality like “workflow” in these platforms can also help manually build workflows, but they have been specifically optimized for integration with LLMs.
Shortcuts app (formerly Workflow(Apple, 2024a) is an app developed by Apple for building workflows through a graphical interface, available on Apple’s operating systems (iOS/iPadOS and macOS). Shortcuts app can be seen as the DAP of Apple. It allows users to create workflows (known as shortcuts (Apple., 2024)) that execute specific tasks on their devices and share them online via iCloud (Apple, 2024c). Users can also download curated shortcuts from the Gallery of the Shortcuts app. However, the shortcuts available in the Gallery are very limited, with only a few dozen options. To access more shortcuts, users must either collect them from third-party sharing sites like Shortcuts Gallery (Gallery, 2024) and SSPAI (SSPai, 2024) or create their own. Shortcuts can be triggered through the Shortcuts app, widgets, the share sheet, and Siri, and they can also be automated to run upon specific events.
Shortcuts are composed of multiple API calls (actions). An agent can use the shortcut as a whole API or utilize the individual APIs involved in the shortcut. This paper treats the various APIs within the shortcuts as APIs available to the agent, aiming for the agent to automatically construct workflows of API calls.
3 Dataset
In this Section, we first introduce the acquisition of the dataset (Section 3.1). Then, we outline the ShortcutsBench’s construction process (Section 3.2). Finally, we outline the setup for evaluation tasks to evaluate the agent’s ability to handle tasks of varying difficulty (Section 3.3.1), including the ability to select suitable APIs, the ability to do parameter filling (Section 3.3.2), and the awareness in requesting additional information from the system or user (Section 3.3.3).
3.1 Dataset Acquisition
Figure 1: (a) illustrates the data acquisition process. (b) shows the dataset acquisition of existing work (Table 1). APIs in existing benchmarks/datasets are created by hand or modified from existing datasets, or API hubs, and queries and action sequences are constructed using templates or semi/fully automated methods with LLMs.
Figure 1 shows the data acquisition process. We first use search engines to identify popular public shortcut-sharing sites ①. We totally find $14$ sites such as Shortcuts Gallery (Gallery, 2024) and SSPai (SSPai, 2024). Then we crawled these sites to obtain fields such as “shortcut name”, “function description”, “shortcut type”, and “iCloud link” ②. After deduplicating based on “iCloud link” (Apple, 2024c), we got the source files of all $8675$ shortcuts (Apple, 2024b) ③. Subsequently, we extracted “app name” using the  field WFWorkflowActionIdentifier in the source file like com.apple.Music, and then downloaded related apps ④ from various sources: (1) third-party apps from the “macOS App Store” or the official website of the app, (2) non-system apps (uninstallable) like Keynote from path /Applications/ on macOS, (3) system apps like Reminders (uninstallable) from path /System/Application/ on macOS, (4) iOS apps from the “iOS App Store”, (5) Shortcuts app itself from path /System/Library/PrivateFrameworks/WorkflowKit.framework/ on macOS. During the downloading, we also excluded some legacy apps and $12$ paid apps. For more details about the whole acquisition process, please refer to Appendix.
Then we managed to extract APIs from the downloaded apps ⑤. The APIs are mainly from intent definition file \${filename}.actionsdata from AppIntent (Inc., 2024b; 2022) framework, \${filename}.intentdefinition from SiriKit (Inc., 2024c; 2022) framework, and WFActions.json from system path /System/Library/PrivateFrameworks/WorkflowKit.framework/ on macOS. We extracted all APIs involved in the app’s shortcuts. During the extraction, due to the existence of multi-platform versions of most apps like Day One (Day One, 2024a; b), and the presence of similar and duplicate API definition files from the same and different frameworks (Inc., 2024c; b; 2022), we perform deduplication based on manually crafted rules. For more details about deduplication, please refer to Appendix. Finally, as shown in Table 1, we get $88$ apps from various categories such as “Health & Fitness”(iTunes App Store, 2024b), “Developer Tools” (iTunes App Store, 2024a), and “Lifestyle” (iTunes App Store, 2024c). These apps in total include $1414$ APIs, including all of $556$ APIs involved in $7628$ shortcuts.
3.2 Dataset Construction
As shown in Figure 2, existing benchmarks/datasets consist of two parts: (1) APIs; (2) queries and corresponding action sequences.
APIs
(“a” in Figure 2) include the “API description” (“a.4”), “API name” (“a.1”), “parameter names” (“a.2”), “parameter types” (“a.1”), “default value” (“a.2”), “return value type” (“a.3”), and “return value name” (“a.3”). The field names in [...] in Figure 2 correspond to the field names in \${filename}.actionsdata, \${filename}.intentdefinition, and WFActions.json. For details, please refer to the Appendix. In existing benchmarks/datasets (Table 1), the “parameter types” (“a.1”) and “return value types” (“a.3”) are composed of primitive data types such as int and string. In addition to primitive data types, APIs in ShortcutsBench also include “enum” or “advanced data types”. Enum is composed of “the class name” and “the possible value”, with each value equipping a “value name”. We also provide the agent with a description of the “enum” in the API information. Advanced data types, such as the model (“a.1”) in app chatGPT (OpenAI, 2024), include three String types named identifier, title, and subtitle. We can comprehend them through their “type name” and “type description”.
Figure 2: The construction of ShortcutsBench. (a) shows the information of API com.openai.chat.AskIntent extracted from the app ChatGPT’s \${filename}.actionsdata. We provide this API description to the LLM, expecting it to call the API at the appropriate time. The API information shown in (a) includes the API functionality description (a.k. (a.1)~(a.4)) as shown in (d), and the user-friendly natural language description of the API (a.k. (a.4.3)) seen by shortcut developers during programming, as shown in (c). (e) presents the shortcut name and functionality description from the shortcut sharing-sites. (b) shows the simplified prompt fed to GPT-4o, instructing it to generating queries based on demands indicated by shortcuts by integrating the info from (c), (d), and (e). Different colors indicate different information sources.
Query and action sequence.
A query is a user command, such as “Tell me what the weather will be like tomorrow.” The action sequence (aka. shortcut) is the series of API calls to complete the query, with each API call referred to as an action. The action sequence identifies the steps needed to complete a query. As shown in Figure 1.b, Existing benchmarks/datasets (Table 1) collect APIs first and then use them, either fully automatically or semi-automatically, to construct query and action sequences through LLMs. In contrast, action sequences in ShortcutsBench are all human-annotated (①②③ in Figure 1). The shortcut developers are our annotators. APIs in ShortcutsBench (④ in Figure 1) are also all real-world.
Generating queries. As shown in Figure 1, existing works construct query and action sequences based on available APIs. In contrast, we construct queries based on existing action sequences and APIs. When constructing a query for a specific action sequence, we need to understand the functional description of the action sequence (“e” in Figure 2) and detailed information about the involved APIs (“a” in Figure 2). With this information, we can generate higher-quality queries. To ensure the quality of the generated queries, we also leverage the unique advantage of shortcuts: the natural language workflow descriptions (“b.7” / “c” in Figure 2). By inputting these intuitive natural language descriptions into an LLM, we can generate more accurate queries. When generating queries, we also require the model to naturally include primitive data type parameters and enum data type needed for API calls in generated queries. This helps us evaluate the agent’s ability to fill in primitive parameters in Section 3.3.2. To ensure the quality of generated queries, we use the state-of-the-art LLM, GPT-4o (OpenAI, 2024b), to generate the queries. The prompt templates we used to generate queries can be found in the Appendix.
3.3 Task Definition and Metrics
We aim to address $3$ research questions regarding the performance of existing agents built using leading LLMs on ShortcutsBench with varying difficulties: (1) How do they perform in API selection? (2) How do they handle API parameter value filling, including parameters for primitive data types, enums, and outputs from previous actions? (3) Can they recognize when input is required for tasks that need system or user information?
Table 2: Final evaluation set with varying difficulties.
$\bm{|aseq_{i}|}$ (0, 1] (1, 5] (5,15] (15,30] Overall
# Queries 706 2169 1571 774 5220
# Avg APIs 1.76 3.94 8.57 13.78 6.50
# Avg Acts 1.00 3.15 9.43 21.58 7.48
Preliminaries. ShortcutsBench consists of a set of queries $Q=\{q_{1},q_{2},...,q_{n}\}$, corresponding "golden" action sequences $ASeq=\{aseq_{1},aseq_{2},...,aseq_{n}\}$, and all available APIs $APIs=\{api_{1},api_{2},...,api_{m}\}$. For each query $q_{i},1\leq i\leq n$, the corresponding “golden” action sequence is $aseq_{i}=\{a_{1},a_{2},...,a_{\left|aseq_{i}\right|}\}$, where the length of the action sequence is $\left|aseq_{i}\right|$. Each app $app_{j}$ has a set of APIs $apis_{j}=\{api_{1},api_{2},...,api_{\left|apis_{j}\right|}\}$. The action sequences generated by the agent for each query $q_{i}$ are referred to as $bseq_{i}$.
Prepare available APIs for each query. For each query $q_{i}$, we provide the LLM with a certain number of usable APIs to simulate real-world scenarios where APIs can be input into the LLM’s context. Following existing work (Meta, 2024; Qin et al., 2023; Tang et al., 2023; Xu et al., 2023; Schick et al., 2024; Hao et al., 2024), we equip each $q_{i}$ with a specific number of APIs. For each $aseq_{i}$, let $\left|APIs_{i}\right|$ represent the number of APIs involved. In addition to these $\left|APIs_{i}\right|$ APIs, we equip each query with extra APIs calculated as $max(min(x\times\left|APIs_{i}\right|,20-\left|APIs_{i}\right|),0)$, where $x\in\{3,4,5\}$. We do this because it is impractical to input all APIs into the context simultaneously. When dealing with a large number of APIs, additional retrieval is often required (Qin et al., 2023), which we do not consider in this work.
Further Processing. Considering the context limitations of LLMs, we excluded shortcuts longer than $30$ and parts using the API is.workflow.actions.runworkflow to call other shortcuts. While these shortcuts remain in our open-source dataset, they will not be included in the subsequent evaluation. We aim to study the performance of agents on queries of varying difficulties. As shown in Table 2, we categorize ShortcutsBench into $4$ difficulty levels and $8$ task types based on $\left|aseq_{i}\right|$ and “shortcut type” (Section 3.1), respectively. For more details, please refer to the Appendix. When calculating the length, for branching actions like is.workflow.actions.conditional, we consider the longest branch as the length. Additionally, we ignore the lengths of looping actions like is.workflow.actions.repeat.count and special actions such as is.workflow.actions.comment. Due to the presence of branching actions, the average number of APIs involved when $p=1$ is greater than one, specifically $1.76$. For a detailed process, please refer to the Appendix. The number of shortcuts in each level is denoted as $n_{p}$. Each query and action sequence is referred to as $q_{p,i}$ and $aseq_{p,i}$, with $1\leq p\leq 4$ and $1\leq i\leq n_{p}$.
3.3.1 Performance about API Selection
Following existing work (Huang et al., 2023; Patil et al., 2023; Li et al., 2023b; Xu et al., 2023; Schick et al., 2024; Hao et al., 2024), we use the accuracy of API selection as the metric. The accuracy is calculated as the number of correct API selections $m_{p}$ divided by $n_{p}$. Specifically, each time we predict an action $b_{j},1\leq j\leq|aseq_{i}|$, we provide the agent with all the correct historical actions $\{a_{1},a_{2},...,a_{j-1}\}$. We then require the agent to predict the next action. All actions predicted by the agent form the prediction sequence $bseq_{p,i}$. This method is similar to the next token prediction (NTP) in LLMs, effectively preventing a cascade of errors in subsequent action predictions due to a single incorrect prediction. During the prediction, when encountering special actions such as branching and looping, we skip predicting these actions and directly add them to the historical actions. For more details, please refer to Appendix. We chose API selection accuracy over the final result for the following two additional reasons:
  • ShortcutsBench contains numerous APIs such as opening the “All Shortcuts Folder” in the Shortcuts app that do not have a return value. This makes it challenging to evaluate using existing metrics that measure the success rate of solving queries (Qin et al., 2023; Tang et al., 2023; Xu et al., 2023; Schick et al., 2024; Xu et al., 2023; Hao et al., 2024).
  • ShortcutsBench includes numerous APIs with complex input and output types, such as PDFs and Rich Text. Converting these formats into text that an LLM can process presents a significant challenge (Naveed et al., 2023), as LLMs struggle to serialize them into text. Consequently, it becomes difficult to ascertain the correctness of the final results. However, measuring API selection accuracy is straightforward.
3.3.2 Effectiveness of API Parameter Value Filling
In this part, we aim to investigate the performance of agents in API parameter value filling, including parameters for “primitive data types” and “enums” and filling output from previous actions. For each input parameter of every action in ShortcutsBench, we expect the agent to fill in the following parameters correctly:
  • Static Parameters Preset: These are static parameters that users provide as default inputs of the action. These static parameters typically include primitive data types such as String and Integer, as well as custom Enum defined by app developers. When the query explicitly specifies a parameter that can be used as a static parameter, we expect the agent to accurately fill in the parameter values according to the user’s query and the API’s definition. When generating queries, we have already required the LLM to naturally include primitive and enumerated data types (Section 3.2). To further ensure that the corresponding parameters are indeed included in the queries during evaluation, we used the LLM to filter these parameters further, ensuring their presence in the queries. Detailed prompts can be found in the Appendix.
  • Outputs from Previous Actions: An action may either have no output or, if it does have an output, the output may be used by the following actions. In shortcuts, In ShortcutsBench, outputs that are difficult to input directly into the LLM are represented by a unique identifier (UID) and an output name (OutputName), which can be input into the LLM for processing. The agent should have the ability to correctly use the output values of previous actions.
For the static parameters preset, we evaluate using the overall parameter fill rate. Let $sppa_{i}$ be the total number of parameters that need to be filled in $aseq_{i}$, $1\leq i\leq n_{q}$, where $n_{q}$ is the number of queries. If the agent correctly fills $sppt_{i}$ parameters in the generated action sequence $bseq_{i}$, then the static parameter preset accuracy can be calculated as $Acc_{spp}=\sum_{i=1}^{n_{q}}sppt_{i}/\sum_{i=1}^{n_{q}}sppa_{i}$. Similarly, for outputs from previous actions, the accuracy can be calculated as $Acc_{ofpa}=\sum_{i=1}^{n_{q}}ofpat_{i}/\sum_{i=1}^{n_{q}}ofpaa_{i}$.
3.3.3 Recognition of Need for Input
In this section, we aim to investigate the ability of existing API-based agents to ask systems or users for necessary information to resolve the missing information. This missing information can come from the system like clipboard (Clipboard), input files (ExtensionInput), and the current date (CurrentDate) or from the user (Ask(Inc., 2024a). For example, a parameter named tags is usually represented in a shortcut as "tags":{"Value":{"Type": "Ask"}}, where "Type": "Ask" indicates that the parameter will prompt the user for input. For a detailed process, please refer to Appendix. We use the proportion of correctly identified parameters to evaluate the agent’s ability to recognize the need for input from the system or the user. Let $n_{s}$ be the number of queries, $aska_{i}$ be the number of times the need from the system or the user appears in $aseq_{i}$, $askt_{i}$ be the number of times the need from the system or the user appears in $bseq_{i}$. The accuracy of ask for necessary information can be calculated as $Acc_{afni}=askt_{i}/aska_{i}$.
4 Evaluation
4.1 Setup
Model. Referencing existing work (Huang et al., 2023; Qin et al., 2023; Patil et al., 2023; Tang et al., 2023; Li et al., 2023b; Xu et al., 2023; Zhuang et al., 2024; Schick et al., 2024; Hao et al., 2024), considering the performance of mainstream LLMs, we selected and tested $9$ most advanced LLMs to construct API-based agent. The chosen model including $4$ closed-sourced LLMs like Gemini-1.5-Pro (DeepMind, 2024), Gemini-1.5-Flash (DeepMind, 2024), GPT-3.5-turbo (OpenAI, 2024b), and ChatGLM-4-Air (ChatGLM, 2024), and $5$ open-source LLMs like LLaMA-3-70B (Meta, 2024), QWen-2-70B (Qwen, 2024b), QWen-2-57B Qwen (2024a), Deepseek-2-Chat (236B) (DeepSeek, 2024b), and Deepseek-2-coder (236B) (DeepSeek, 2024a). Among them, Gemini-1.5-Pro, LLaMA-3-70B, QWen-2-70B, Deepseek-2-chat, and Deepseek-2-coder are LLMs benchmarked against GPT-4 OpenAI (2024a), while Gemini-1.5-Flash, ChatGLM-4-Air, and QWen-2-57B are benchmarked against GPT-3.5-turbo performance. We did not evaluate smaller LLMs like LLaMA-3-8B (Hugging Face, 2024) or Vicuna-7b-v1.5 (LMSYS, 2024) because we found that agents built on them can only handle simple tasks such as single API selection (Huang et al., 2023) and they cannot handle well on advanced tasks like parameter filling. Agents built with such models often fail to produce the required JSON actions correctly and frequently generate nonsensical outputs.
Prompt Template. Following existing work (Huang et al., 2023; Qin et al., 2023; Patil et al., 2023; Tang et al., 2023; Li et al., 2023b; Xu et al., 2023; Zhuang et al., 2024; Schick et al., 2024; Hao et al., 2024), we slightly modified the ReACT (Yao et al., 2022) templates to construct the API-based agents. Specifically, we added prompts related to shortcuts, such as the types of fillable parameters and the meanings of special statements like branches. For all $3$ research questions (RQs), we use the same prompt templates. An agent should correctly select APIs, fill in parameters, and be aware of the need to request necessary information from the system or user at appropriate times. Please refer to Appendix for more details.
4.2 Result Analysis
Figure 3: The API selection accuracy on queries with different complexity levels.
Figure 4: The API selection accuracy difference of each LLM across $8$ task types.
Figure 5: The API selection accuracy of each task type on $9$ API-based agents.
Through the results of API selection accuracy (Section 3.3.1), we get the following conclusions:
  • Agents built using open-source LLMs now perform comparably to closed-source models on lower-difficulty tasks but still lag on higher-difficulty tasks. From Figure 4 we know that open-source LLMs $>=70$B match the performance of closed-source LLMs from the first $3$ difficulty tasks, significantly outperforming GPT-3.5-turbo. However, they still lag behind closed-source LLMs in handling complex tasks at the $4$-th level. Moreover, the price of open-source LLMs is significantly lower than that of GPT-3.5-turbo. For more details, please refer to Appendix.
  • Existing LLM-based agents still perform poorly on tasks requiring multi-step reasoning, even Gemini-1.5-Pro level LLMs struggle with high-difficulty tasks. From Figure 4 we know that almost all LLMs handle well in API selection tasks at the level of (0,1], but only more advanced models like Gemini-1.5-Pro and QWen-2-72B can do well in higher-difficulty tasks of (1,5]. As tasks become more complex, the accuracy drops sharply. The average accuracy dropped by $19\%$ as task difficulty rose from (0,1] to (1,5], ranging from a $9\%$ decrease (Deepseek-2-chat) to a $44\%$ (ChatGLM-4-Air). From (0,1] to (5,15], accuracy fell by $46\%$, with drops from $38\%$ (Gemini-1.5-Pro) to $58\%$ (ChatGLM-4-Air).
  • Agents built with the same LLM show significant performance variations across different types of tasks. From Figure 5 we know that the performance difference of agents built with different LLM ranges from $18.04$% (ChatGLM-4-Air) to $36.70$% (Gemini-1.5-Pro).
  • Existing API-based agents perform well on tasks in daily life such as Lifestyle & Social but show poorer performance on professional tasks like Development & API. From Figure 5 we know that Lifestyle & Social exhibit the highest average accuracy, surpassing the lowest category, Development & API by approximately $18\%$.
Based on the results of API Parameter Value Filling (Section 3.3.2), we draw following conclusions:
(a) Accuracy of primitive data types & enum data types (upper) and outputs from previous actions (lower).
(b) The error rates for action parameter value filling.
  • Compared to the API selection, for existing most intelligent LLM like Gemini-1.5-Pro, increased task difficulty has a much smaller impact on the accuracy of parameter filling, especially on using outputs from previous actions. As shown in Figure 6(a), the precision of API parameter filling of the existing most intelligent LLM like Gemini-1.5-Pro and QWen-2-72B remains similar across tasks of varying difficulty in both the upper and lower figures. This indicates that the greatest limitation of existing API-based agents in addressing user queries lies in the reasoning and planning capabilities implied by API selection.
  • Compared to API selection, the performance of API parameter filling remains a bottleneck for existing cost-effective LLMs like GPT-3.5-turbo and ChatGLM-4-Air. As shown in Figure 6(a), the performance of these LLMs in API parameter filling significantly decreases as task difficulty increases.
  • Compared to using the outputs of previous actions, extracting relevant parameters from the user’s query and filling them according to the query and API description is more challenging. As shown in Figure 6(a), the colors in the top plot (filling primitive data types and enum data types) are generally lighter than those in the bottom plot (filling the outputs of previous actions as parameters). The accuracy drop ranges from $2.55$% (GPT-3.5-turbo) to $15.39$% (Deepseek-2-Chat).
  • For existing cost-effective LLMs like GPT-3.5-turbo and ChatGLM-4-Air, errors mainly stem from incorrect output formats and wrong API selections. Figure 6(b) shows error types for tasks requiring outputs from previous actions. It can be seen that powerful LLMs like  Gemini-1.5-Pro rarely make format errors, whereas the most cost-effective models frequently make mistakes in both output format and API selection.
The results from Recognition of Need for Input (Section 3.3.3) lead us to the following conclusions:
  • All agents perform poorly at recognizing necessary system and user inputs when required. Overall, all agents have weak recognition capabilities, with accuracy ranging between $30.55$% (GPT-3.5-turbo) and 55.18%(Deepspeed-2-coder). Larger LLMs such as Deepspeed-2-chat (236B) still demonstrate better recognition accuracy.
Table 3: The accuracy of recognition of the need for input from the system or the user.
Levels Gemini QWen Deeps. Deeps. LLaMA Gemini QWen GPT ChatGLM
1.5-Pro 2-72B 2-chat 2-coder 3-70B 1.5-Flash 2-57B 3.5-turbo 4-Air
(0, 1] 33.33 37.78 64.29 62.71 47.62 62.79 22.22 28.89 47.62
(1, 5] 45.95 50.40 55.50 60.08 44.08 53.99 37.24 37.70 48.06
(5, 15] 51.85 36.42 40.76 49.44 35.71 40.65 28.37 20.33 48.42
(15, 30] 46.67 25.00 27.59 43.14 22.22 44.64 8.11 17.14 48.89
Overall 46.59 41.97 47.90 55.18 49.89 40.71 30.74 30.55 48.28
5 Conclusion
In this paper, we introduce ShortcutsBench, a benchmark for evaluating API-based agents. To the best of our knowledge, ShortcutsBench is the most realistic, rich, and comprehensive benchmark of its kind. Our findings indicate that for agents built on the most advanced LLMs, the primary bottleneck is API selection. For the most cost-effective LLMs, there is considerable room for improvement in both API selection and parameter filling. Additionally, we identified a significant deficiency in the agents’ awareness of requesting necessary information.